Question 8 3 pts Like Victoria from the previous questions, David has $2000 to put toward consumption this month. Also like Victoria, David believes there is a 40% chance he will have a bike accident this month, in which case he expects to incur medical costs of $1500 (leaving him with $500 to put toward consumption). Unlike Victoria, though, Davids utility over consumption is given by Suppose again that the premium for $1500 in insurance coverage is $750, and that the premium for $500 in insurance coverage is $250 (i.e., neither is priced at an actuarially fair rate). Calculate Davids expected utility associated with having no insurance as well as with each of the two insurance options (full and partial). Based on your new answers, fill in each of the blanks below for the case in which insurance is not actuarially fair with>, , or- EUlno insurance) EU(partial insurance) EUlno insurance) EUlfull insurance) EU(partial insurance) EUlfull insurance) In each blank, enter only >, <, or =. Each entry is worth 1 point. Hint: In this question, it is important to pay attention to the decimal places (at least out to the hundredth place).
Answer #1

If there is no insurance cover, then the cost of accident is $1500

Then there will be $500 left for consumption and if there is no accident then there will be $2000 for consumption.

Expected utility = 5000.2*0.4 + 20000.2 * 0.6 = 4.13012

If there is a partial insurance of $500 then the premium paid will be $250 and now the cost of accident will be $1000. So the total cost is $1250

So now only $750 will be left for consumption. And if there is no accident amount available for consumption will be $1750

Expected utility = 7500.2*0.4 + 17500.2*0.6 = 4.17491

If there is full insurance, then the premium paid will be $750. The amount left for consumption will be $1250. We will not be considering the probability of accident here because we will not be paying any medical cost if accident occurs.

So the expected utility will be = 12500.2*1 = 4.16277

So with the values calculated we can say that ,

EU(no insurance) < EU( partial insurance)

EU(no insurance) < EU(full insurance)

EU(partial insurance) > EU(full insurance)